# ViWikiFC: Fact-Checking for Vietnamese Wikipedia-Based Textual Knowledge Source

### v1

@@ Abstract
Fact-checking is essential due to the explosion of misinformation in the media ecosystem. Although false information exists in every language and country, most research to solve the problem mainly concentrated on huge communities like English and Chinese. Low-resource languages like Vietnamese are necessary to explore corpora and models for fact verification. To bridge this gap, we construct ViWikiFC, the first manual annotated open-domain corpus for Vietnamese Wikipedia Fact Checking more than 20K claims generated by converting evidence sentences extracted from Wikipedia articles. We analyze our corpus through many linguistic aspects, from the new dependency rate, the new n-gram rate, and the new word rate. We conducted various experiments for Vietnamese fact-checking, including evidence retrieval and verdict prediction. BM25 and InfoXLMLarge achieved the best results in two tasks, with BM25 achieving an accuracy of 88.30% for SUPPORTS, 86.93% for REFUTES, and only 56.67% for the NEI label in the evidence retrieval task, InfoXLMLarge achieved an F1 score of 86.51%. Furthermore, we also conducted a pipeline approach, which only achieved a strict accuracy of 67.00% when using InfoXLMLarge and BM25. These results demonstrate that our dataset is challenging for the Vietnamese language model in fact-checking tasks.
keywords:
@@
Fact Checking, Language Model, Information Verification, Corpus
1 Introduction
Following the increase in the amount of information, the lack of strict policies when spreading information has led to misinformation and disinformation on social media. This can cause conflict and manipulate the group of people. In an effort to minimize the impact of misinformation and disinformation in society, some organizations like PolityFact [^1] https://www.politifact.com/, FactCheck.org [^2] https://www.factcheck.org/, play an essential role in online Fact Verification when manually verifying claims based on a different source of evidence. Manual verification is a time-consuming task since a Fact-Checker would have to search through many potential sources for the evidence relevant to the claim. This work is insufficient with the speed of updated information on social media.
Automatic Fact-Checking is a complicated task. In a recent survey on fact-checking [1] that can be separated into four sub-tasks: claim detection, where a machine has to choose what claim needs to be verified; evidence retrieval is to find evidence that SUPPORTS or REFUTES the claim, and verdict prediction is an attempt to determine the veracity of the claim based on evidence in the evidence retrieval task. Finally, the justification production is conducted with a reasonable explanation of why a claim is supported or refuted by the evidence. However, there are many works on how to improve the performance of fact-checking systems in different ways, from applying fast dot product indexing in evidence retrieval [2, 3], constructing new reasoning methods for the verification task [4] to a large-scale corpus with natural input such as NELA [5], FakeCovid [6], to artificial input for research purposes such as FEVER [7], VitaminC [8], languages as Vietnamese are still considered a low resource for natural language processing (NLP) due to lack of corpora, especially in automatic fact-checking.
To contribute to the process of NLP research for Vietnamese, especially in Fact-Checking, we present the first large-scale, open-domain corpus for Vietnamese Fact-Checking on Wikipedia, ViWikiFC: Vietnamese Wikipedia Fact-Checking. The corpus consists of 20,916 claims manually annotated and based on evidence retrieved from Wikipedia pages. Our corpus takes advantage of FEVER [7]; our corpora also have three label classes: SUPPORTS, REFUTES, and NOTENOUGHINFORMATION (NEI). In our corpora, the evidence is manually rewritten into three types of claims based on three class labels. In this way, claims are more realistic when in reality and very diverse in many semantic ways from the way claim construction in the FEVER corpus [7] based on just one piece of evidence. Furthermore, we also follow ViNLI [9] corpus creation to create a double claim with different meanings for each label from one evidence to capture more than just one piece of information in evidence.
Before corpus creation, we developed a suitable guideline and annotation tool for the annotation process. Annotators who are Vietnamese native speakers with an education background beyond high school were carefully trained with the guidelines and a strict training process to ensure corpus creation consistency.
We conducted two experiments, evidence retrieval, and verdict prediction, to evaluate the challenge of ViWikiFC. For the evidence retrieval task, the input is a claim sentence and a document corpus. We use TF-IDF and BM25 for lexical match, and Vietnamese-SBERT [10] for semantic search to retrieve a sentence most relevant to the claim from the document corpus. For the verdict prediction (VP), we used three deep neural networks, including CBOW [11], BiLSTM [12]. We also evaluated using SOTA pre-trained language models such as Multilingual BERT (mBERT) [13], XLM-R [14], and InfoXLM [15], as well as monolingual models designed for Vietnamese, such as PhoBERT [16], ViT5 [17], and ViDeBERTa [18]. These models have shown exceptional ability in performing the VP task. Ultimately, we have seamlessly integrated the top-performing models for both tasks into a comprehensive pipeline and thoroughly assessed the effectiveness of the resulting system on a test set.
Contributions in this paper are described as follows.
  • Firstly, we propose the first fact-checking corpus on Vietnamese Wikipedia, comprising 20,916 claim sentences manually annotated based on 73 Wikipedia articles built with a consensus agreement among annotators, achieving a 95.87% Fleiss’$\kappa$-agreement.
  • Next, we analyze the experimental results of language models from different aspects of the corpus to highlight the challenges introduced by our corpus.
  • Finally, our corpus and annotation tool is available for research purposes [^3] https://github.com/HighWill0/ViWikiFC.git.
Our paper is structured into five sections: Section 1 provides an overview of the problem and our contributions, Section 2 introduces related research on fact-checking, including existing corpora (see Section 2.1) and baselines (see Section 2.2). Section 3 describes the corpus construction process, including annotator recruitment and training (see Section 3.1), annotation and validation tool (see Section 3.2, selection of data sources (see Section 3.3), claim generation (see Section 3.4), corpus validation (see Section 3.5), and analysis of the linguistic aspects of the corpus (see Section 3.6). Section 4 introduces the baseline models tested on various tasks (see Section 4.1), summarizes the experimental results (see Section 4.2), and analyzes the impact of the data set on the methods (see Section 4.3). Finally, in Section 5, we conclude our research findings and propose future directions for the fact-checking task in Vietnamese.
2 Related Work
In this section, we introduce the corpus in Section 2.1 as well as the baseline models in Section 4.1 for the fact-checking task.
2.1 Fact-Checking Corpora
We performed a review of existing fact-checking datasets, as summarized in Table 1. We categorized the datasets based on six attributes: data set size (claims), language, data source, domain, annotated evidence, and annotated label
@@ 2.1 Fact-Checking Corpora
The emergence of the Politifact corpus [19] constructed by Vlachos and Riedel in 2014 laid the foundation for investigating the verification of claims with a structure made up of three main tasks, known as claim detection, evidence retrieval, and verdict prediction, in recent fact-checking studies. However, the corpus still carried the limitation of being too small, restricting the training and evaluation of neural network-based models and following a similar approach to the Politifact corpus [19], Liar [20] was published with more than 12.8K samples extracted from the Politifact [^4] https://www.politifact.com/ website. However, the accuracy prediction in this corpus relied solely on claims or metadata without utilizing external sources and did not provide any evidence for verification.
Fever [7] introduced a large-scale corpus with more than 180k claim sentences manually written by a team of annotators based on Wikipedia [^5] https://www.Wikipedia’s.org/ pages. This corpus presented an automated information verification process by combining models for distinct tasks into a complete pipeline. In subsequent years, new corpora were introduced, such as Snopes [21] with more than 6K samples and MultiFC [22] with more than 36k samples collected from various fact-checking websites such as PolitiFact [^6] https://www.politifact.com/. Particularly in the MultiFC [22] corpus, claims were collected from multiple sources, resulting in label counts ranging from 2 to 27. From 2020 onward, a series of corpora appeared with various approaches. For example, Hover [23] had more than 26k samples and was built and verified using multiple Wikipedia pages. The VitaminC corpus [8] was built with more than 488k samples based on actual revisions of Wikipedia, all including evidence.
Apart from the Corpora constructed from regular text data, there are also Corpora for semi-structured data, such as tables. TabFact [24], with more than 92k samples, and InfoTabs [25], with more than 23k samples, were constructed based on tables from Wikipedia. The FEVEROUSE corpus [26], with more than 87K samples, uses both text and a table for verification. Corpora in other languages have been developed and researched as well. X-Fact [27] with more than 31K samples in 25 different languages excluding English, or ANT [28] with more than 4K samples for Arabic, DanFEVER [29] with more than 6K samples for Danish, and the most recent CHEF corpus [30] with 10K samples for Chinese.
With the growing demand for the development of automatic fact-checking tasks, especially in Vietnamese NLP research, an open-domain, large-scale, and high-quality corpus becomes crucial. Following the corpus development efforts, we constructed a Vietnamese Fact-Checking corpus alongside an information verification process using pre-trained language models in the Vietnamese language, contributing to advancing research and addressing Fact-Checking challenges in Vietnamese.
2.2 Fact-Checking Methods
As previously stated, fact-checking encompasses four distinct sub-tasks: claim detection, evidence retrieval, verdict prediction, and justification production. Of these sub-tasks, evidence retrieval and verdict prediction are particularly crucial, as they serve as the basis for determining ground-truth information and information inference mechanisms to verify information. Our research focuses on these two sub-tasks, so in this section, we introduce the methods used in the two sub-tasks.
Evidence retrieval is the task of searching and extracting information from various sources and types of data, such as text, tables, images, and knowledge bases, to verify a claim. In recent research, evidence retrieval is often approached through two main directions. First, the traditional non-neural approach, such as TF-IDF [7, 30] and BM25 [31], which uses sparse representations to measure term matching between a claim and evidence. Although this method doesn’t require much time and resources, it can only capture lexical information and is limited in handling various queries. The second approach is neural network-based, which can calculate and measure the semantic similarity between evidence and a claim. In 2020, Karpukhin et al. introduced dense passage retrieval using a dual-encoder architecture to represent questions and passages as dense vectors. This approach has successfully incorporated BERT-based language models and their variants into open-domain question-answering [32, 33, 34] and fact-checking tasks [35, 36]. Cross-encoder architecture is also utilized for re-ranking passages and sentences by BERT-based and its variants re-rankers [37, 38, 39], which achieved significant improvement compared to traditional methods.
Verdict prediction is a task to verify claims based on evidence extracted from the evidence retrieval task. The input for this task consists of a pair of sentences, including evidence and claims. However, for datasets with natural claims [5, 22, 6] or datasets built for research purposes [24, 23], the claim needs to be inferred and verified from information in multiple pieces of evidence. In recent years, solutions have been proposed, such as concatenating all pieces of evidence into one [40] or using special components like kernel-based attention [41] or a graph-based approach [42]. These approaches create conditions for verifying information on more complex claims [43, 44, 45]. Early approaches for this task used neural network-based methods such as CBOW [11], RNN [46], and BiLSTM [12] to verify claims, achieving high performance. In 2018, BERT [47], a pre-trained language model based on the transformer architecture [48], significantly improved in various natural language processing tasks, including verdict prediction. In the following years, many pre-trained language models have been developed, such as RoBERTa [49], ALBERT [50], and XLM-R [14], which are multilingual pre-trained language models. The first monolingual pre-trained language model for Vietnamese is PhoBERT [16] developed with the same architecture as BERT [47] and the same pre-training technique as RoBERTa [49]. PhoBERT was trained on a 20GB word-level Vietnamese Wikipedia and News corpus, achieving state-of-the-art results in many downstream tasks such as NER, POS, and NLI. Since then, many pre-trained language models for Vietnamese have been developed, such as ViT5 [17], a transformer-based encoder-decoder model with the same architecture and training method as T5 [51], and ViDeBERTa [18], which uses the DeBERTa architecture [52].”
3 Corpus Creation
@@
3.1 Annotator Recruitment and Training
We hired thirty annotators who are Vietnamese native speakers and have high school to college literacy to create claims from the evidence given. These annotators are in the habit of reading and searching for information on the Internet. Annotators must undergo a strict training process with two phrases before entering the official annotation process; we use the Fleiss $\kappa$ score to calculate the agreement of annotators with the authors.
In the first phase, the authors trained the annotators through the provided guidelines to familiarize themselves with the rules when creating claims. Next, the annotators were required to create a claim in the training corpus that contained 50 evidence sentences. The annotator training corpus labels (SUPPORTS, REFUTES, and NOTENOUGHINFORMATION) were marked. One of Three authors read and verdict the label for every claim-evidence pair; if the proportion of the labels agreed upon by Two other authors is over 0.90, the annotator is selected for the second phase of the training process. If the result is not over 0.95, the annotator must learn the mistake from the previous training corpus and go through the first training process again with another training corpus.
In the second phase, thirty annotators are randomly divided into six distinct groups, each consisting of five individuals. Before this phase, the thirty individuals worked on at least one data folder containing 150 claim-evidence pairs. The thirty data folders have their labels concealed. Each group is assigned five data folders (already label-hidden) from any of the remaining five groups. All members within a group then relabel the data folders of the members in the other designated group. Labeling results are used to calculate the level of consensus using the Fleiss $\kappa$ score [53]. If the consensus level among the five annotators on a data folder is greater than or equal to 0.95, the Annotator who labels that folder is selected for the official annotation process. In cases where the consensus level is below 0.95, the Annotator must go through the first training process. The procedure for calculating consensus is repeated for all annotators in different groups until all annotators have been evaluated for consensus. During test annotation, we encourage annotators to list exceptional cases (multi-meaning in a sentence, unknown subject) and note frequent mistakes in the guideline.
3.2 Fact-Checking Annotation Tool
In order to create the environment for corpus creation, we developed a manual labeling tool for annotators. The tool is built using the Java Development Kit  [^7] https://www.oracle.com/java/, which makes it compatible with any operating system and easy to use. Figure 2 shows a complete screen after finishing the generated claim in Section 3.4. During claim generation, the annotator must follow some strict rules, including the following.
  • Rule 1: Generating all six claims in one sentence of evidence before moving on to the next.
  • Rule 2: The annotator must accurately write the claim when generating the claim.
  • Rule 3: The first letter of a claim must be capitalized, and the end must be followed by a period.
  • Rule 4: Regional or specific dialect words should be avoided in writing a claim.
Furthermore, we also built a validation tool (see Figure 3) in the Corpus Validation (see Section 3.5).
Figure 2: ViWifiFC annotation tool for fact-checking corpus.
Figure 3: ViWifiFC validation tool for fact-checking corpus.
3.3 Evidence Selection
@@
3.4 Claim Generation
Rather than creating only two labels representing SUPPORTS or REFUTES, we add a NOT ENOUGH INFORMATION label as in FEVER [7]. In reality, we encounter claims that do not have enough evidence to prove that they are entirely leaning toward a side. We ask annotators to create claim sentences for the following three labels.
  • SUPPORTS: Annotators create a claim that we can determine to be true based only on information from the evidence.
  • REFUTES: Annotators create a claim that we can determine is false based only on the information in evidence.
@@ 3.4 Claim Generation
  • NOT ENOUGH INFORMATION (NEI): Annotators create a claim that we cannot determine to be true or false based only on the evidence provided.
For two labels, SUPPORTS and REFUTES, the annotator composes a claim sentence based on the information provided by the evidence sentence. When writing the claim, the annotator is not allowed to use personal understanding or external knowledge beyond the evidence. When crafting the claim sentence, the annotator is free to use any information provided by the evidence sentence and to adopt various writing styles to improve the complexity and diversity of expression in the claim that is different from the generation of FEVER claims [7] when a single piece of information conducts the claim. For NOTNOUGHINFORMATION labels, the annotator uses at least one subject, object, or event in the evidence sentence, plus some information outside of it but still in the context of the evidence sentence. In this way, we can construct an evidence-retrieval experiment for this label in the section.
Annotators are required to write two distinct claim sentences for each corresponding label for each evidence sentence, resulting in six claim sentences for the three labels from a single evidence sentence. The construction of claim sentences is analogous to the approach used in the two corpora: OCNLI [54] and IndoNLI [55], where the annotators are instructed to take advantage of multiple pieces of information provided by the evidence sentence and use various expressions for the same information template. Furthermore, for a given evidence sentence, the claim sentences corresponding to different labels exhibit differentiation in writing style and the information used to compose the claim.
During the Annotator training process, we observed that a significant portion of annotators encountered difficulties in composing claim sentences with the SUPPORTS label, especially for evidence sentences with limited exploitable information, without excessively mirroring the evidence sentence in the Paraphrase. This occurs with evidence sentences that have a length of 15 to 25 tokens. Additionally, the REFUTE label poses challenges for annotators when constructing sentences by negating the information in the evidence sentence. However, frequent and dense negation may lead to a shift from the REFUTE label to the NEI label. Table 3 illustrates REFUTED claims with too many information changes; if not written and carefully reviewed, they tend to lean towards the NEI label.
Sentence
Evidence Hitler gây ra sự kiện ”Đêm của những con dao dài”, giết hại các đối thủ của ông. (Hitler caused the ”Night of the Long Knives” event, killing his opponents.)
Claim Hitler chịu trách nhiệm cho việc giết hại các đối thủ trong sự kiện ”Đêm của những con dao dài”. (Hitler was responsible for killing his opponents during the ”Night of the Long Knives”.)
Table 2: Example of generating SUPPORTS claim based on evidence with few tokens.
Sentence
Evidence Con đường tơ lụa được nhà Tùy chú trọng kế thừa và phát triển vào thời kỳ hưng thịnh của triều đại này vào thế kỷ thứ 7. (The Silk Road was emphasized for inheritance and development by the Sui dynasty during its prosperous period in the 7th century.)
Claim Con đường tơ lụa vẫn được kế thừa và xây dựng dù nhà Đường bước vào giai đoạn suy vong của thế kỷ thứ 10. (The Silk Road continued to be inherited and developed even as the Tang dynasty entered a period of decline in the 10th century.)
Table 3: Example of generating REFUTED claims using more than two rules lead to NEI label.
3.5 Corpus Validation
In the corpus evaluation process, we establish a two-part evaluation procedure, namely annotator validation and author validation. Both of these procedures are conducted during the claim generation phase to ensure a high level of consensus throughout the entire labeling process.
3.5.1 Annotator Validation
During the claim generation phase, upon reaching 50.00% corpus completion, we identify the top 5 annotators who have made significant contributions to the data construction. Subsequently, we select a distinct corpus these 5 Annotators have not previously labeled. This new corpus comprises 750 pairs of claim-evidence sentences. The existing set of labels on this corpus is then removed. All five annotators collaboratively apply labels to each sentence pair until all 750 claim-evidence pairs have been re-evaluated. In cases where fewer than three of the five annotators assign an identical label to a sentence pair, that particular sentence is designated with a ”-” mark, and those pairs are subsequently eliminated from the corpus. The agreement rate among the annotators reached 99.87% when at least three annotators were assigned the same label, and the agreement rate when at least annotators were assigned the same label was 95.47%. Additionally, we measured the inter-annotator agreement using Fleiss $\kappa$ score, and the achieved result was 95.87%, higher than the agreement rate in the FEVER corpus, which reached only 68.41%.
3.5.2 Author Validation
We used 750 pairs of claim-evidence sentences, which were relabeled in Section 3.5.1 and evaluated whether the written claim sentences adhere to the guidelines. We observed that 96.12% of claim sentences followed the stipulated guidelines, while 3.87% of the claim sentences contained writing errors and did not adhere to the annotation guideline rules.
3.6 Corpus Analysis
3.6.1 Overall Statistics

### v2

@@ Abstract
Fact-checking is essential due to the explosion of misinformation in the media ecosystem. Although false information exists in every language and country, most research to solve the problem has mainly concentrated on huge communities like English and Chinese. Low-resource languages like Vietnamese are necessary to explore corpora and models for fact verification. To bridge this gap, we construct ViWikiFC, the first manually annotated open-domain corpus for Vietnamese Wikipedia Fact Checking more than 20K claims generated by converting evidence sentences extracted from Wikipedia articles. We analyze our corpus through many linguistic aspects, from the new dependency rate, the new n-gram rate, and the new word rate. We conducted various experiments for Vietnamese fact-checking, including evidence retrieval and verdict prediction. BM25 and InfoXLMLarge achieved the best results in two tasks, with BM25 achieving an accuracy of 88.30% for SUPPORTS, 86.93% for REFUTES, and only 56.67% for the NEI label in the evidence retrieval task. InfoXLMLarge achieved an F1 score of 86.51%. Furthermore, we also conducted a pipeline approach, which only achieved a strict accuracy of 67.00% when using InfoXLMLarge and BM25. These results demonstrate that our dataset is challenging for the Vietnamese language model in fact-checking tasks.
@@
Keywords: Fact Checking, Language Model, Information Verification, Corpus
[…]
ViWikiFC: Fact-Checking for Vietnamese Wikipedia-Based Textual Knowledge Source
Hung Tuan Le, Long Truong To, Manh Trong Nguyen, Kiet Van Nguyen
University of Information Technology, Ho Chi Minh City, Vietnam
Vietnam National University, Ho Chi Minh City, Vietnam
{21520250, 21521101, 21520343}@gm.uit.edu.vn, kietnv@uit.edu.vn
Abstract content
1.  Introduction
11footnotetext: Corresponding author.
@@ 1.  Introduction
Following the increase in the amount of information, the lack of strict policies when spreading information has led to misinformation and disinformation on social media. This can cause conflict and manipulate the group of people. To minimize these impacts, organizations like PolitiFact and FactCheck.org play an essential role in online Fact Verification by manually verifying claims based on different sources of evidence. However, manual verification is time-consuming and insufficient given the speed of information updates on social media.
Automatic Fact-Checking is a complicated task that can be separated into four sub-tasks: claim detection, evidence retrieval, verdict prediction, and justification production Guo et al. (2022). While there are many works on improving fact-checking systems Lewis et al. (2020); Maillard et al. (2023); Pan et al. (2023) and large-scale corpora for various languages such as NELA Horne et al. (2018), FakeCovid Shahi and Nandini (2020), FEVER Thorne et al. (2018), and VitaminC Schuster et al. (2021), Vietnamese remains a low-resource language for NLP, especially in automatic fact-checking.
To contribute to Vietnamese NLP research, we present ViWikiFC: the first large-scale, open-domain corpus for Vietnamese Fact-Checking on Wikipedia. The corpus consists of 20,916 manually annotated claims based on evidence from Wikipedia pages. Following FEVER Thorne et al. (2018), our corpus has three label classes: SUPPORTS, REFUTES, and NOTENOUGHINFORMATION (NEI). The evidence is manually rewritten into three types of claims, making them more realistic and semantically diverse than FEVER’s approach. We also follow ViNLI Huynh et al. (2022) to create double claims with different meanings for each label from one evidence.
Annotators who are native Vietnamese speakers with education beyond high school were carefully trained to ensure consistency. We conducted experiments on evidence retrieval using TF-IDF, BM25, and Vietnamese-SBERT Phan et al. (2022b), and on verdict prediction using CBOW Mikolov et al. (2013), BiLSTM Graves and Schmidhuber (2005), multilingual models (mBERT Kenton and Toutanova (2019), XLM-R Conneau et al. (2020), InfoXLM Chi et al. (2021)), and Vietnamese-specific models (PhoBERT Nguyen and Tuan Nguyen (2020), ViDeBERTa Tran et al. (2023)).
Contributions in this paper are described as follows.
  • Firstly, we propose the first fact-checking corpus on Vietnamese Wikipedia, comprising 20,916 claim sentences manually annotated based on 73 Wikipedia articles built with a consensus agreement among annotators, achieving a 95.87% Fleiss’k-agreement.
  • Finally, our corpus and annotation tool are available for research purposes.
2.  Related Work
In this section, we review prior studies and resources related to fact-checking, focusing on corpora and baseline models that have shaped this research direction. Section 2.1 introduces key datasets, while Section 2.2 presents the main methodological approaches.
2.1.  Fact-Checking Corpora
We reviewed existing fact-checking datasets, summarized in Table 1, categorized by dataset size, language, source, domain, and annotation types.
The PolitiFact corpus Vlachos and Riedel (2014) laid the foundation for modern fact-checking by defining three core tasks: claim detection, evidence retrieval, and verdict prediction. However, its limited size restricted model training. The LIAR dataset Wang (2017) expanded on this with over 12.8K samples from PolitiFact [^1] https://www.politifact.com/ but lacked evidence annotations. FEVER Thorne et al. (2018) later introduced a large-scale corpus of 180K manually written claims based on Wikipedia [^2] https://www.wikipedia.org/, establishing an automated pipeline structure for verification. Subsequent corpora such as Snopes Hanselowski et al. (2019) and MultiFC Augenstein et al. (2019) diversified sources and label types, while later datasets like HoVer Jiang et al. (2020) and VitaminC Schuster et al. (2021) expanded evidence coverage and reasoning complexity.
For semi-structured data, TabFact Chen et al. (2020), InfoTabs Gupta et al. (2020), and FEVEROUS Aly et al. (2021) introduced fact verification using tables or hybrid text-table formats. Multilingual efforts have also emerged, including X-Fact Gupta and Srikumar (2021), ANT Khouja (2020), DanFEVER Nørregaard and Derczynski (2021), and CHEF Hu et al. (2022).
Despite this progress, Vietnamese NLP still lacks an open-domain, large-scale, and well-annotated fact-checking corpus. To address this gap, we constructed a Vietnamese fact-checking dataset and developed an accompanying verification pipeline using pre-trained Vietnamese language models.
@@ 2.1.  Fact-Checking Corpora
$2.$  Fact-Checking Methods
Fact-checking typically involves four sub-tasks: claim detection, evidence retrieval, verdict prediction, and justification generation. Among these, evidence retrieval and verdict prediction are central, forming the core of our research.
Evidence retrieval involves finding relevant information from various sources to support or refute a claim. Traditional non-neural methods such as TF-IDF Thorne et al. (2018); Hu et al. (2022) and BM25 Chen et al. (2017) rely on lexical matching but are limited in semantic understanding. Neural approaches, such as Dense Passage Retrieval (DPR) Karpukhin et al. (2020), use dense vector representations via dual-encoder architectures, achieving significant improvements in open-domain QA Karpukhin et al. (2020); Chang et al. (2019); Lee et al. (2019) and fact-checking Samarinas et al. (2021, 2020). Cross-encoder re-rankers based on BERT and its variants Nogueira and Cho (2019); Nogueira et al. (2019); Wang et al. (2019) further enhance retrieval precision.
Verdict prediction determines the truthfulness of a claim based on retrieved evidence. It is typically formulated as a sentence-pair classification problem. Early approaches used CBOW Mikolov et al. (2013), RNN Elman (1990), or BiLSTM Graves and Schmidhuber (2005), while transformer-based models such as BERT Devlin et al. (2019) and its successors—RoBERTa Liu et al. (2019), ALBERT Lan et al. (2019), and XLM-R Conneau et al. (2020)—greatly improved performance. In Vietnamese, PhoBERT Nguyen and Tuan Nguyen (2020) was the first monolingual pre-trained model, trained on a 20GB corpus of Wikipedia and news data, followed by ViT5 Phan et al. (2022a) and ViDeBERTa Tran et al. (2023), which further enhanced representation quality for downstream tasks, including fact-checking.
3.  Corpus Creation
@@
3.1.  Annotator Recruitment and Training
We recruited thirty Vietnamese native speakers with at least high school to college literacy to create claims based on given evidence sentences. All annotators were familiar with online reading and information searching. Before participating in the official annotation, they underwent a two-phase training process to ensure annotation quality, with agreement measured by the Fleiss $\kappa$ score Fleiss (1971).
In the first phase, annotators were trained using detailed guidelines to understand claim-generation rules. They then annotated a small training corpus containing 50 evidence sentences labeled as SUPPORTS, REFUTES, or NOT ENOUGH INFORMATION. Each claim-evidence pair was reviewed by one of three authors, and inter-author agreement above 0.90 was required for the annotator to proceed to the second phase. Annotators scoring below 0.90 were asked to review their mistakes and repeat the first phase using a new training set.
In the second phase, the thirty annotators were randomly divided into six groups of five. Each annotator labeled at least one dataset folder of 150 claim-evidence pairs, with labels concealed. Every group then re-labeled five folders produced by other groups, and inter-annotator agreement was again evaluated using Fleiss $\kappa$. Annotators achieving a consensus score of 0.95 or higher qualified for the official annotation process; otherwise, they repeated the first training phase. This procedure continued until all annotators achieved the required agreement threshold. During pilot annotations, annotators were encouraged to document ambiguous cases (e.g., polysemy, unclear entities) and note common labeling errors to refine the annotation guidelines.
3.2.  Evidence Selection
@@
$3.$  Claim Generation
Following the FEVER dataset Thorne et al. (2018), we design three labels—SUPPORTS, REFUTES, and NOT ENOUGH INFORMATION (NEI)—to capture the different relationships between a claim and its supporting evidence. In practice, many claims cannot be conclusively verified or falsified due to incomplete or ambiguous information; therefore, we include the NEI label to better reflect these real-world scenarios. Annotators are instructed to compose claim sentences for each of the three labels as follows:
  • SUPPORTS: The claim can be determined as true based solely on the given evidence.
  • REFUTES: The claim can be determined as false based solely on the given evidence.
@@
  • NOT ENOUGH INFORMATION (NEI): The claim cannot be judged as true or false using the provided evidence.
For the SUPPORTS and REFUTES labels, annotators write claims that rely only on the content of the evidence sentence, without using external knowledge or personal understanding. They are encouraged to reformulate, restructure, or vary the expression to enhance linguistic diversity and avoid simply paraphrasing the evidence. This approach differs from FEVER Thorne et al. (2018), where claims often derive from minimal factual transformations.
For the NEI label, annotators must still reuse at least one component from the evidence (subject, object, or event) while introducing new but contextually relevant information not present in the original sentence. This design ensures the NEI label remains semantically related to the evidence and also enables a more realistic evaluation of evidence retrieval and claim verification models.
Each evidence sentence leads to the creation of two distinct claim sentences for each label, resulting in a total of six claims per evidence. This procedure follows the approach in OCNLI Hu et al. (2020) and IndoNLI Mahendra et al. (2021), where annotators are instructed to explore multiple information pieces within a single evidence and use diverse linguistic forms to represent similar meanings. Furthermore, for a given evidence sentence, the claims corresponding to different labels show clear variation in writing style, word choice, and focus, thereby improving the dataset’s semantic diversity.
During annotator training, we observed that a noticeable portion of annotators struggled with constructing SUPPORTS claims, especially when the evidence sentence contained limited information (typically 15–25 tokens). It was challenging for them to compose a natural claim without overly mirroring the evidence. Similarly, the REFUTES label posed difficulties due to excessive or unclear negation, which occasionally caused label drift toward NEI.
3.4.  Corpus Validation
To ensure data quality and annotation consistency, we perform a two-stage validation process: annotator validation and author validation, both conducted during the claim generation phase.
3.4.1.  Annotator Validation
When 50% of the corpus was completed, the top five annotators were selected to relabel a separate subset of 750 claim–evidence pairs that none had previously worked on. Labels were re-assigned collaboratively, and pairs with fewer than three matching labels were discarded. The inter-annotator agreement reached 99.87% when at least three annotators agreed and 95.47% when at least two agreed. The Fleiss $\kappa$ score was 95.87%, significantly higher than FEVER’s 68.41%, confirming strong consistency among annotators.
3.4.2.  Author Validation
The same 750 pairs were further reviewed by the authors to assess guideline compliance. Results show that 96.12% of the claims strictly followed the annotation rules, while only 3.87% contained minor writing or labeling errors, demonstrating the overall reliability of the corpus.
3.5.  Corpus Analysis
3.5.1.  Overall Statistics